### v1

# A geometric phase approach to quark confinement from stochastic gauge-geometry flows

Abstract
We apply a stochastic version of the geometric (Ricci) flow, complemented with the stochastic flow of the gauge Yang–Mills sector, in order to seed the chromo-magnetic and chromo-electric vortices that source the area-law for QCD confinement. The area-law is the key signature of quark confinement in Yang–Mills gauge theories with a non-trivial center symmetry. In particular, chromo-magnetic vortices enclosed within the chromo-electric Wilson loops instantiate the area-law asymptotic behaviour of the Wilson loop vacuum expectation values. The stochastic gauge-geometry flow is responsible for the topology changes that induce the appearance of the vortices. When vortices vanish, due to topology changes in the manifolds associated to the hadronic ground states, the evaluation of the Wilson loop yields a dependence on the length of the path, hence reproducing the perimeter law of the hadronic (Higgs) phase of real QCD. Confinement, instead, is naturally achieved within this context as a by-product of the topology change of the manifold over which the dynamics of the Yang–Mills fields is defined. It is then provided by the Aharonov–Bohm effect induced by the concatenation of the compact chromo-electric and chromo-magnetic fluxes originated by the topology changes. The stochastic gauge-geometry flow naturally accomplishes a treatment of the emergence of the vortices and the generation of turbulence effects. Braiding and knotting, resulting from topology changes, namely stochastic fluctuations of the Einstein–Yang–Mills system, stabilise the chromo-magnetic vortices and dynamically induce, as non-trivial topological features, chiral symmetry-breaking. The latter is due to the appearance of torsional degrees of freedom, to which four fermion interactions are related, and to the formation of fermion condensates, in which gluon strings get stabilised by non-trivial braiding.
I Introduction
The underlying mechanism of confinement in quantum Yang–Mills theories, with the related understanding of the way quarks in Quantum Chromo Dynamics (QCD) are bound into hadrons, has been a hot topic of investigation for decades. The community is currently seeking to reach a deeper understanding beyond semiclassical explanations that involve the possible crucial role of chromo-electric strings and of magnetic monopoles’ condensates, as developed in seminal papers by Mandelstam, ’t Hooft and Polyakov Mandelstam (1976); Zichichi (1975); Polyakov (1975) (for a recent review of the underlined concepts and implications, see e.g. Refs. Greensite (2011, 2017a); Pasechnik and Šumbera (2021) and references therein).
An insight about the relevance of topology for confinement was offered by ’t Hooft in his renown paper ’t Hooft (1978). Sets of operators that modify the gauge-topological structure of gauge field theories were considered, showing that physical effects nonetheless can be local. In $2+1$ dimensions, it was demonstrated that in absence of a spontaneous breakdown of the local gauge symmetry, topological fields acquire a vacuum expectation value (vev) that breaks spontaneously their mutual global symmetry. This mechanism then induces permanent confinement of quarks. In $3+1$ dimensional non-abelian gauge theories it was then shown that topological operators and gauge field operators close an algebra that explains the occurrence of four possible phases: i) spontaneous breakdown induced by a Higgs field; ii) permanent confinement of gauge quantum numbers in the absence of the Higgs field, either explicit or composite; iii) confinement in presence of the Higgs mechanism; iv) an intermediate phase, presumably a critical point, holding massless particles. The phases i) and ii) were then shown, for a simple realisation of the algebra, to be connected by dual transformations. Phase ii) was called superinsulating, as dual to the super-conductive phase. Along these lines, Diamantini, Trautenberger and Vinokur proposed in Diamantini et al. (2018) that the superinsulating state, characterised by an infinite resistance, may emerge on the insulating side of the superconductor-insulator transition in superconducting films, providing a realisation of confinement with direct experimental access.
Topology entered hitherto without involving geometric considerations, but solely referring to the properties of the vacuum states, intended as background states. Here, we propose a programme towards a dynamical description of confinement in QCD that is rooted in geometric topology, geometry, and stochastic quantisation techniques. The theoretical framework we develop intertwines among concepts of topology and out-of-equilibrium physics. In our framework, the quark confinement phenomenon is described as a geometric-phase effect that emerges dynamically, according to a stochastic flux along thermal time that drives the system toward relaxation. The stochastic flux implements a scale transformation in the Renormalisation Group (RG) sense, and is responsible for the change of topology of the manifold associated to the QCD ground-state under scrutiny. Nucleons, as well as mesons and more complex composite resonances, are indeed seen as low-curvature manifolds of increasingly high topology classes characterised by the stochastic flow.
The change of topology in the manifolds, on the other hand, is interpreted in the dual picture as the generation of flux-lines of the gauge connection (gluon field). These flux-lines source Wilson loops and center vortices ultimately reproducing the asymptotic area law of the Wilson-loop vevs. The phase transition between the perimeter- and the area-law regimes shall therefore be addressed in terms of the topology changes that are induced by the stochastic geometric flow. This latter is coupled to the stochastic RG flow of the gauge field through the out-of-equilibrium dynamics in the thermal time. The stochastic gauge-geometry flux then drives the relaxation of the system toward equilibrium configurations, these latter being characterized by the Einstein–Yang-Mills equations.
The programme we propose involves the following three stages:
  • i) Establishing the framework of topological RG (TRG), a version of the Ricci flow (RF) tailored to topological quantum field theories (TQFTs). While the RF implements rescaling of topologies through conformal transformations, TRG complements the RG with changes of topology;
  • ii) Advocating the stochastic quantisation method to leverage RF and TRG. The stochastic geometric flow, or stochastic RF (SRF) is then extended to the Wilsonian formulation of Yang–Mills theories, considering their background interaction with the gravitational field, affecting the topology of the electric and magnetic Yang–Mills background configurations;
  • iii) Developing a geometric phase approach to confinement. The geometric curvature is not relevant to the purpose of these studies. What really matters is topology: the reason to consider the interaction with gravity is therefore finding the dynamics that induces through changes of topology, the generation of vortices in the Yang–Mills gauge sector, which finally source the area-law scaling.
Achieving these steps realises a novel proposal for a dynamical realisation of QCD confinement, a non-perturbative phenomenon responsible for a plethora of quark-gluon and hadron physics implications in the infrared regime of QCD.
The plan of the paper is the following. In Sect. II we provide an overview of the center vortex mechanism of confinement, serving as a baseline for further considerations. In Sect. III we outline the (stochastic) gauge-geometry flow formulation and its implication for the dynamics of confinement. In Sect. IV we elaborate on the important details of the out-of-equilibrium dynamics in he thermal time and on the stochastic RF. The formulation of the Wilsonian approach to QCD and dynamical stochastic evolution of topology in the context of confinement is provided in Sect. V. Then, in Sect. VI we present a novel picture of QCD confinement dynamics resorting to the geometric (topology) fluctuations due to the stochastic gauge-geometry flow. Conclusions and outlook follow in Sect. VII. Finally, in Appendix A we provide details of $3$-manifold topology and geometry relevant for our formulation.
II Confinement and center vortices
In this section, we summarise the basics of the vortex mechanism of QCD confinement Engelhardt et al. (1998a) that features a number of numerical evidences as summarised in Ref. Greensite (2017a). A detailed review of this and other mechanisms of confinement is beyond the purpose of this paper and can be found e.g. in Refs. Greensite (2011); Pasechnik and Šumbera (2021) and references therein. Therefore, here we only provide a brief introduction into the underlined concepts relevant for our subsequent formulation, pointing to the relevant literature where further details can be found.
Center symmetry
The Elitzur’s theorem Elitzur (1975) ensures that the phases of a gauge theory cannot be distinguished in terms of the breakdown of local gauge symmetries. Therefore, the breakdown of additional global symmetries must be considered that enables to identify gauge phases also in presence of a local order parameter — this is indeed the case of the $\mathbb{Z}_{2}$ symmetry in the Ising model. At this point, we remind that fixing a covariant gauge connection does not fix completely the gauge freedom, leaving instead remnant symmetries that evade the Elitzur’s theorem and may therefore undergo spontaneous symmetry breaking. This is the case of the Kugo-Ojima condition Kugo (1995); Kugo and Ojima (1979), a suitable confinement criterion that requires the full residual symmetry in the Landau gauge $\partial^{\mu}A_{\mu}^{a}=0$ to remain unbroken, hence implying the vanishing in any physical state $\psi$ of the expectation value of $\langle\psi|Q_{a}|\psi\rangle$, the color charge operator. The global and space-time dependent parts of this full residual gauge symmetry with respect to the gauge transformation $A_{\mu}\rightarrow GA_{\mu}G^{\dagger}$ in the Landau gauge acquires the form Hata (1982, 1983)
$G(x)$ $=$ $\exp\left(\frac{\imath}{2}\Xi^{a}(x)\sigma^{a}\right)\,,$ (1)
$\Xi^{a}(x)$ $=$ $\epsilon^{a}_{\mu}x^{\mu}-g\frac{1}{\partial^{2}}(A_{\mu}\times\epsilon^{\mu})^{a}+\mathcal{O}(g^{2})\,,$ (2)
with $\epsilon^{a}_{\mu}$ arbitrary parameters and $g$ the $SU(2)$ gauge coupling. In addition to (1), also a further criterion must be fulfilled for confinement: that another part of the residual gauge symmetry that is space-time independent remains unbroken.
The separation of phases is characterised by non-analytic boundaries between magnetic order and disorder configurations, which are associated with (either spontaneous or dynamical in the thermal time) breaking of global center symmetries in Yang–Mills gauge groups. Specifically, a center symmetry is a subset of gauge group elements that commute with all the elements of the gauge group — e.g. $\mathbb{Z}_{N}$ for $SU(N)$, where the group elements are $\exp 2\pi\imath n/N$, with $n=0,\dots N-1$. The representations of $\mathbb{Z}_{N}$ individuate $N$ possible subsets, known as $N$-alities, belonging to infinite numbers of possible representations of $SU(N)$. A given $N$-ality $k$ corresponds to the number of boxes of Young tableau mod $N$. Therefore, $N$-alities represent how a given representation transforms under a given center symmetry group. For instance, for $R$ representations of $SU(N)$ group elements $g$, i.e. $R(g)$, the action of center symmetries is expressed as $R(zg)=z^{k}\,R(g)$, with $z$ being an element of $\mathbb{Z}_{N}$ and $k$ the $N$-ality class. The Yang–Mills action
$\mathcal{S}_{\rm W}=-\frac{1}{4g^{2}}{\rm Tr}[W_{\circ\gamma}(A)+W_{\circ\gamma}^{\dagger}(A)]$ (3)
is written in terms of the super-trace of the Wilson loops $W_{\circ\gamma}(A)$. These are holonomy operators $H_{\gamma}(A)$, depending on the gauge-field $A$, that are traced with respect to the representation indices and supported on loops $\circ\gamma$, the closure of the curves $\gamma$.
The singular gauge transformation, holding implicit dependence on the thermal time, applied to $H_{\gamma}(A)$ reads
$H_{\gamma(\tau)}(A)\rightarrow G_{\gamma(\tau_{s})}H_{\gamma(\tau_{0})}(A)G_{\gamma(\tau_{t})}^{\dagger}\,,$ (4)
where $\tau_{s,t}$ are the thermal times at the source ($s$) and the target ($t$), and the periodicity condition
$G_{\circ\gamma}=z^{*}G_{\circ\gamma}$ (5)
holds up to a center symmetry transformation. A particular case of this transformation applied to $H_{\gamma}(A)$ is
$H_{\gamma(\tau_{0})}(A)\rightarrow z\,H_{\gamma(\tau_{0})}(A)\,,\quad{\rm with}\quad z\in\mathbb{Z}_{N}\,,$ (6)
with $\tau=\tau_{0}$ selecting a fixed thermal time-slice. This transformation, in the asymptotic thermal time limit, corresponds to
$A_{\mu}(x)\rightarrow G(x)A_{\mu}(x)G^{\dagger}(x)-\frac{\imath}{g}G(x)\partial_{\mu}G^{\dagger}(x)\,.$ (7)
The transformation (7) realises at equilibrium an “almost” gauge symmetry, but only for $\mu=0$, and only at the base point of the loop, where the second term (turned into a $\delta$-function) is dropped — hence, the notion of a “singular” transformation. Such a transformation does not leave the action invariant, but has a special role to be elaborated upon more in what follows.
Unbroken center symmetry in the pure Yang–Mills theory hence provides an important criterion for confinement, which is associated with an area-law fall-off of large Wilson loop vevs, effectively indicating the emergence of a magnetic disorder state. In our picture, which will be detailed in what follows, the out-of-equilibrium dynamics induces the relaxation of the system from magnetic order to magnetic disorder configurations, hence from the deconfined phase to a confined one. Indeed, a magnetically disordered configuration implies a linear growth of the potential of static quarks asymptotically with the distance between them, thus yielding a linear string potential.
As the action appears to be non-invariant under singular gauge transformations, consequently, a loop of magnetic flux with a singularity is effectively induced by such a transformation. Furthermore, we conjecture that a singular gauge transformation is tightly connected to the out-of-equilibrium dynamics, being associated with the creation of thin center vortices in the gauge-field system.
Geometric phase
A well-known example is provided by the abelian $U(1)$ case, whose closed holonomy on a spacelike loop $\circ\gamma$ provides a geometric phase
$W_{\circ\gamma}=e^{\imath e\Phi_{B}}\,,$ (8)
with $\Phi_{B}$ magnetic flux through $\circ\gamma$ and $e$ electric charge. When the loops winds around a solenoid oriented along the $z$-axis, one may find configurations corresponding to non-vanishing magnetic fluxes even for zero magnetic fields along $\circ\gamma$. This latter can be viewed as a result of the singular gauge transformation applied to a vanishing gauge field with a discontinuous $G(x)$. In cylindrical coordinates $\{r,\theta,z,t\}$ this entails a discontinuity in $\theta$ for $r>0$, namely,
$W_{\circ\gamma}\rightarrow e^{\pm\imath e\Phi_{B}}W_{\circ\gamma}\,,$ (9)
with $\exp{\pm\imath e\Phi_{B}}$ a $U(1)$ group element and $\pm$ depending on the orientation of $\circ\gamma$ and corresponding to the production along the $z$-axis of a singular line of magnetic flux. The winding number is defined as the number of revolutions around a fixed point in two space dimensions. This can be also generalised in three space dimensions through the so-called linking number, specifying how many times the two loops wind around each other.
For a $SU(N)$ Yang–Mills theory in $d$ dimensions the geometric phase is replaced by a center-group element of $\mathbb{Z}_{N}$, namely
$G_{\gamma(\tau_{s})}=zG_{\gamma(\tau_{t})}\,,\qquad W_{\circ\gamma}\rightarrow(z^{*})^{l}W_{\circ\gamma}\,,$ (10)
with $l$ linking number and $\circ\gamma$ a space-like Wilson-loop, topologically linked to $(d-2)$-dimensional thin vortices. The quanta of these non-abelian magnetic fluxes are referred to as thin center vortices. The regularisation of the singular color-magnetic field is then achieved by smearing out the vortices in directions transverse to the two-dimensional hypersurfaces. These are known as thick center vortices Greensite (2011).
’t Hooft loop
The ’t Hooft loop can be introduced as the operator $B_{\circ\gamma}$ that creates thin center vortices at the base point $x_{0}=x(\tau_{0})$ of a loop $\alpha\equiv\circ\gamma$, embedded in the base manifold $M$ of the principal bundle of the gauge theory. If two loops $\alpha$ and $\alpha^{\prime}$ on a three-dimensional hypersurface are topologically linked, with linking number $l=1$, then by our previous discussions the following relation holds:
$B_{\alpha}W_{\alpha^{\prime}}=zW_{\alpha^{\prime}}B_{\alpha}\,,\qquad z\in\mathbb{Z}_{N}\,.$ (11)
The vev of a Wilson loop $W_{\alpha}$, which we denote as ${\cal W}_{\alpha}\equiv\langle W_{\alpha}\rangle$, and of a ’t Hooft loop, namely, $\langle B_{\alpha}\rangle$ in their asymptotic behaviour, either satisfy a perimeter-law or an area-law, but never simultaneously, as shown in ’t Hooft (1978).
For $a$ and $b$ positive and real numbers, a confined (magnetically disordered) phase is achieved when
$\mathcal{W}_{\alpha}\sim e^{-aA_{\alpha}}\quad\leftrightarrow\quad\langle B_{\alpha}\rangle\sim e^{-bP_{\alpha}}\,,$ (12)
while a deconfined phase is enforced, for $a^{\prime}$ and $b^{\prime}$ positive and real numbers, when
$\mathcal{W}_{\alpha}\sim e^{-a^{\prime}P_{\alpha}}\quad\leftrightarrow\quad\langle B_{\alpha}\rangle\sim e^{-b^{\prime}A_{\alpha}}\,.$ (13)
The latter case implies a spontaneously broken center symmetry, entailing a magnetically ordered phase. Therefore, the Wilson loops and the ’t Hooft loops can be considered as dual to one another: the former creates loops of color electric fluxes, while the latter creates closed loops of color-magnetic flux (thin center vortices), both at fixed thermal time $\tau$.
Wilson loops and characterisation of confinement
The Wilson loops operators, as closed holonomies, are represented as the path-ordered (typically, denoted as “$\mathcal{P}$”) exponential of the gauge connection $A_{\mu}$, namely,
[…]
The path contour that is customarily picked is a rectangular time-like loop that describes the creation, the propagation and finally the annihilation of a pair of static quark and anti-quark that are placed at space-like base-points of the loop. In the Euclidean, this can be obtained, for instance, by combining operators that create particle-antiparticle pairs in a color singlet state at a given time $T_{E}$ and separation $R$, i.e.
$\mathcal{C}(T_{E})=\psi^{\dagger}(0,T_{E})H_{\gamma_{R}}\psi(R,T_{E})\,,$ (15)
where the holonomy of the gauge connection $H_{\gamma_{R}}$ along the line $\gamma_{R}$, from 00 to $R$ at a fixed time $T_{E}$, creates a color-electric flux tube (string) among the charges. Neglecting the mass term of particles in the heavy-mass limit, the action for the gauge-matter system in the Euclidean lattice formulation is customarily written as
$\mathcal{S}=$displaystyle-- $\frac{\beta}{N}\sum_{p}{\rm Tr}\left[U_{\alpha}(A)+\mathrm{c.c.}\right]$ (16)
$+$ $\delta\sum\limits_{x,\mu}\left[\psi^{\dagger}(x)H_{\gamma_{\mu}}(x)\psi(x+\hat{\mu})+\mathrm{c.c.}\right]\,,$
with $\beta$ and $\delta$ lattice couplings, $\hat{\mu}$ denoting a generic direction in the Euclidean space, $H_{\gamma_{\mu}}(x)$ representing the holonomy along paths $\gamma_{\mu}$, oriented along generic directions $\mu$, which close a generic plaquette $p$ around a base point $x$. In the heavy matter fields limit an infinite energy would be required to pull pairs of particles and antiparticles away from the vacuum, placing them on mass-shell. A similar amount of energy would be also required to bind pairs to sources and screen their charges, hence stretching the flux tube (of strings) to an infinite length before they break apart, an ad absurdum argument that supports the picture of confinement of the quantum vacuum.
Integrating out the matter fields $\psi$ in the functional integral while considering the vev of the product operator $\mathcal{C}(T_{E})^{\dagger}\mathcal{C}(0)$ is equivalent, in the heavy mass limit, to considering $\mathcal{W}_{\alpha}(R,T_{E})$, the vev of the Wilson loop, namely
$\langle\mathcal{C}(T_{E})^{\dagger}\mathcal{C}(0)\rangle\sim\mathcal{W}_{\alpha}(R,T_{E})\,,$ (17)
with
$\mathcal{W}_{\alpha}(R,T_{E})=\langle{\rm Tr}[H_{\gamma}H_{\gamma^{\prime}}\dots]_{\alpha}\rangle\equiv\chi_{\alpha}[H(R,T_{E})]\,,$ (18)
where $\chi$ denotes the trace of the representation of gauge group-elements.
In the operator formalism Greensite (2011) one can derive the asymptotic $T_{E}\rightarrow\infty$ limit
$\langle\mathcal{C}(T_{E})^{\dagger}\mathcal{C}(0)\rangle\propto\sum_{n}|c_{n}|^{2}e^{-\Delta E_{n}T_{E}}\sim e^{-\Delta E_{\rm min}\,T_{E}}\,,$ (19)
with $\Delta E_{n}$ the energy of the $n$-th excited states above the vacuum, and $\Delta E_{\rm min}$ the minimum among them, which individuates a dominant term contribution to the vev. The energy difference between two static charges, namely the interaction static potential $V(R)$, corresponds precisely to $\Delta E_{\rm min}$. Therefore, the asymptotic limit of the Wilson loop along the time-like (counterclockwise) rectangular path $\alpha=R\times T_{E}$, is found to be characterised by the potential $V(R)$.
If we now use the Stokes theorem, we can characterise different types of fall-off for the Wilson loops, hence distinguishing among confined and de-confined phases. For the paradigmatic case of $U(1)$, considering a planar and non self-intersecting Wilson loop one can find
[…]
where the integration individuates the magnetic flux over the minimal area enclosed by a large Wilson loop. The additive properties of a loop enables to decompose it into a product of smaller loops, picked in a way that neighbor loops run in opposite directions to one another, i.e.
$W_{\alpha}=\prod\limits_{i=1,n}H_{\gamma_{i}}\,.$ (21)
Magnetic disorder is accomplished when magnetic fluxes through the smaller loops of the decomposition (i.e. plaquette variables) are completely uncorrelated and their vevs factorise, i.e.
$\mathcal{W}_{\alpha}=\prod\limits_{i=1,n}\langle H_{\gamma_{i}}\rangle=\exp\left[-\sigma_{\gamma}A_{\alpha}\right]\,,$ (22)
where
$\sigma_{\gamma}=-\frac{\ln\langle H_{\gamma}\rangle}{A^{\prime}}\,,$ (23)
and $A$ and $A^{\prime}$ are the larger and smaller Wilson loop areas, respectively.
Within the absence of light matter fields, in principle able to screen the color charge of massive sources, and considering a rectangular Wilson loop supported over the path $\alpha=R\times T_{E}$, magnetically disordered states can be shown to be characterised by a linear growth of the interaction potential in the distance $R$ between the asymptotically static charges, namely,
$V(R)=\sigma R+2V_{0}\,,$ (24)
which is the potential of a linear string of tension $\sigma$, in a given representation that does not depend on the sub-loop area $A^{\prime}$, complemented with a self-energy contribution $V_{0}$. The area law is then recovered for $T_{E}\gg R$ for a generic counter that encloses a large minimal area $A_{\alpha}$, i.e.
$\mathcal{W}_{\alpha}\sim\exp\left[-\sigma A_{\alpha}-V_{0}P_{\alpha}\right]\,,$ (25)
with $P_{\alpha}$ denoting the perimeter of the loop $\alpha$ in the limit $T_{E}\gg R$.
In general, in two space-dimensions one can show for any gauge group that only the magnetically disordered phase is achieved and that the area-law fall-off of the Wilson loop is recovered, due to the absence of Bianchi constraints on the components of the strength tensor of the gauge field Halpern (1979). On space hypersurfaces of higher than two dimensions, the Bianchi identities correlate field strength values at neighbor sites, so that they cannot fluctuate independently from one point to another Batrouni and Halpern (1984). The absence of correlation among the arbitrary-size plaquettes in which a Wilson loop is decomposed is indeed at the onset of the area-law relation. On space-hypersurfaces with dimension higher than two, this correlation disappears, enabling the area-law, only in the strong-coupling limit, which corresponds to the leading order in $\beta\ll 1$. Indeed, this cannot occur in the $\beta\gg 1$ (weakly coupled) regime of $3+1$ dimensions electrodynamics, entailing the failure of the area-law and the establishment of a massless phase characterised by the potential Intriligator and Seiberg (1996)
$V(R)=-\frac{{g^{2}(R)}}{R}+2V_{0}\,.$ (26)
This latter corresponds to a perimeter-law fall-off of the Wilson loop vev, i.e.
$\mathcal{W}_{\alpha}\sim\exp[-V_{0}P_{\alpha}]\,,$ (27)
where the coupling constant $g(R)$ becomes a slow function of $R$ and tends to a constant in the Coulomb phase.
In a similar way, the magnetically disordered phase has been derived for non-abelian gauge theories while inspecting large Wilson loops, both by means of the Stokes theorem Arefeva (1980); Fishbane et al. (1981); Diakonov and Petrov (1989); Karp et al. (1999); Hirayama and Ueno (2000); Diakonov and Petrov (2001); Kondo and Taira (2000a, b) and by deploying finite-range correlators behaviours Di Giacomo et al. (2002); Kuzmenko et al. (2004).
Center vortex mechanism for confinement
A simple exposition of the center vortex mechanism of confinement was offered in Engelhardt et al. (1998b), and reviewed in Greensite (2017b). The argument is extremely simple and intuitive. We overview it in this section.
We start considering a planar rectangle of length $L$, which is pierced by $N$ center vortices. Embedded in the plane we consider a closed holonomy along a loop $\alpha$, i.e. a Wilson loop $W_{\alpha}$, enclosing an area $A$. The probability that $n$ among the $N$ vertices will be puncturing the area enclosed by the Wilson loop is provided by the generic factor of the Newton polynomial expansion, i.e.
$P_{N}(n)=\left(\begin{array}[]{c}N\\ n\end{array}\right)\left(\frac{A}{L^{2}}\right)^{n}\left(1-\frac{A}{L^{2}}\right)^{N-n}\,.$ (28)
Piercing of each vortex through the Wilson loop entails a contribution to the vev of the Wilson loop by a factor $(-1)$. Thus, the vev of a Wilson loop can be straightforwardly recovered by weighting the contributions provided by each piercing vortex with the probability distribution for $n$ vortices to puncture the area of the Wilson loop, namely
$\mathcal{W}_{\alpha}=\langle W_{\alpha}\rangle=\sum\limits_{n=0}^{N}(-1)^{n}\,P_{N}(n)=\left(1-\frac{2A}{L^{2}}\right)^{N}\,.$ (29)
Introducing the vortex density $\rho=N/L^{2}$, which one may fix as constant in the infinite volume limit $L\rightarrow\infty$ and $N\rightarrow\infty$, it is immediate to recover the falloff of the area-law:
$\mathcal{W}_{\alpha}=\lim\limits_{N\rightarrow\infty}\left(1-\frac{2\rho A}{N}\right)^{N}=e^{-2\rho A}\,.$ (30)
This mechanism will be at the onset of our formulation, based on out-of-equilibrium stochastic gauge-geometry flow and TQFT.
III Theoretical framework
The theoretical framework our analysis is rooted in comprise topology and out-of-equilibrium physics. Topology is crucial for our discussions for the following reason. We associate to the hadronic background states generic (space-time) manifolds, on which the gauge and quark fields are supported, and the confined configurations are defined. We will argue below that precisely the stochastic (yet, classical) topology transformations instantiate the creation and annihilation of central vortices of the $SU(3)$ connection at the onset of the transitions among confined and deconfined phases in QCD.
On the other hand, creation and annihilation of center vortices, which are understood in terms of stochastic fluctuations in topology, require an out-of-equilibrium approach to the problem. Properties of relaxation toward the infrared hadronic background configurations that are characterised by confinement are derived by considering the out-of-equilibrium version of the Einstein–Yang–Mills equations. The effective action approach, with the calculation of loop-correction to the infrared Lagrangian of non-abelian gauge fields Savvidy (1977), is therefore substituted in this framework by a stochastic flow for the interacting physical systems that comprises the non-abelian gauge fields, the matter (quark) fields and the gravitational field, i.e. geometry. The gravitational field, at the energy scales that are proper of the infrared physics and relevant to unveil confinement in hadronic states, is treated classically. Nonetheless, geometry is still crucial, since it drives the stochastic fluctuations of the underlined topology that enable to create/annihilate center vortices of the gluon fields.
Considering a (stochastic) geometric flow is therefore essential to define the topology of the manifolds associated to hadronic background states. Stochasticity is a necessary requirement for the geometric flow, as it entails breakdown of diffeomorphisms in the infrared/large scale limit, the one at which the gravitational field is measured. Following the seminal paper by Parisi and Wu Parisi and Wu (1981), out-of-equilibrium physics is related to a breakdown of the infrared symmetry that is dynamical in the thermal (“stochastic”) time. The relevance of the Parisi-Wu approach for gravity was further addressed by Rumpf Rumpf (1986), who extended the Parisi–Wu prescription for out-of-equilibrium stochastic quantisation to the Einstein gravity. What is relevant for geometry in our approach is to describe the breakdown of the diffeomorphism symmetry (and the related changes in topology of the manifolds associated to the hadronic systems) of the gravitational field.
On the other hand, the size of these manifolds is still regulated by the dimensional transmutation scales of non-abelian gauge fields. The fermi scale that dictates the effective size of the hadronic system is not affected directly by the stochastic geometric flow.
Stochastic geometric flow
The key idea is based on the fact that in the stochastic geometric (Ricci-like) flow the curvature fluctuations act as a multiplicative noise on the metric field,
$\begin{split}\frac{\mbox{d}}{\mbox{d}\tau}g_{\mu\nu}&=-2\imath\left[R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R\right]+\xi_{\mu\nu}\\ &=-2\imath\mathcal{G}_{\mu\nu\alpha\beta}\frac{\delta S_{G}}{\delta g_{\alpha\beta}}+\xi_{\mu\nu}\,,\end{split}$ (31)
where $\xi_{\mu\nu}=\xi g_{\mu\nu}$ and $\mathcal{G}_{\mu\nu\alpha\beta}$ denotes the super-metric Lulli et al. (2022). The multiplicative noise $\xi_{\mu\nu}$ can be understood as a possible mechanism for multi-fractality, i.e. the non-trivial scaling of the observables cumulants as a function of the scale. This is reflected in a fractal dimension that depends on the scale Lulli et al. (2022).
Given that the tensorial character of the gravitational field is defined against general coordinate transformations, it is convenient to recall the transformation properties of the stochastic time $\tau$, which the reader can find in further details in Lulli et al. (2022). A possible choice is to interpret $\tau$ as the proper time. Its basic property is that it coincides with the coordinate time whenever the metric is almost Minkowskian, as it happens at spatial infinity for the Schwarzschild metric. This renders the equation covariant under general coordinate transformations, given that the proper time is a scalar. As in well-known considerations of turbulence, one can study the fluctuations of the metric components $\delta g_{\mu\nu}(\delta\tau)=g_{\mu\nu}(\tau_{2})-g_{\mu\nu}(\tau_{1})$ at a fixed proper time distance $\delta\tau=\tau_{2}-\tau_{1}$, which sets a characteristic scale.
A necessary step for this analysis is to introduce a coupling to matter fields. This is achieved by considering the “target” Ricci tensor induced by the matter fields
$R^{T}_{\mu\nu}=\frac{8\pi G}{c^{4}}\left[T_{\mu\nu}-\frac{1}{2}g_{\mu\nu}T\right]\,,$ (32)
which enables to rewrite the RF equation as follows:
$\begin{split}\frac{\mbox{d}}{\mbox{d}\tau}g_{\mu\nu}&=-2\imath[R_{\mu\nu}-R^{T}_{\mu\nu}]+\xi_{\mu\nu}\\ &=-2\imath\left[R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R-\frac{8\pi G}{c^{4}}T_{\mu\nu}\right]+\xi_{\mu\nu}\,.\end{split}$ (33)
Here, we have utilised the ansatz about the multiplicative nature of the gravitational stochastic noise, $\xi_{\mu\nu}\equiv\xi g_{\mu\nu}\sim(R-T)g_{\mu\nu}$, and related it to stochastic fluctuations of the trace of the Einstein equation at equilibrium. In stead, the square bracket in the second line of Eq. (33) still corresponds to the first variation of the action describing gravity and its coupling to matter.
The scaling dimension of the cumulants ($n$-points functions) could provide a possible mechanism for a dimensional reduction based on the intermittency of the curvature/matter fluctuations at the small scales Lulli et al. (2022). Knots can arise on average over the large $\delta\tau$ scales from configurations that might not be knotted at smaller scales, just because the effective (fractal) dimension would be different.
Out-of-equilibrium Einstein–Yang–Mills
The Einstein–Yang–Mills (out-of-equilibrium) dynamics, which involves the mutual back-reaction of gauge fields and gravity, is the key to unveil the generation of center vortices. This finally provides the emergence of confinement through configurations supported on non-trivial topologies.
We first introduce the action for the gluon fields which are allowed to fluctuate out of equilibrium in the non-perturbative regime. The gluon fields then acquire a dependence on the thermal time, i.e. $A^{a}_{\mu}(x,\tau)$, while $a$ denotes usual indices of the adjoint representation of the gauge symmetry group. On a curved space-time with Lorentzian signature (,+,+,+)(-,+,+,+)( - , + , + , + ) and metric $g_{\mu\nu}(x)$, the Einstein–Yang–Mills action is expressed by
$\mathcal{S}_{\rm EYM}=$ $+\frac{1}{2\kappa}\int d^{4}x\,\sqrt{-g}\,R[g_{\mu\nu}]$ (34)
$-\frac{1}{4g_{s}^{2}}\int d^{4}x\,\sqrt{-g}\,F^{a}_{\mu\nu}\,F^{a\,\mu\nu}\,.$
For the associated Langevin equations written to describe the out-of-equilibrium dynamics of the system, we implement a multiplicative ansatz for the noise, along the lines of Lulli et al. (2022). The stochastic flow of the gauge field’s components $A^{a}_{\mu}$ is introduced via the stochastic noise $\xi^{a}_{\mu}(x,\tau)\equiv\xi^{\prime}(x,\tau)A^{a}_{\mu}(x,\tau)$. Similarly, for the metric field components the stochastic noise term is denoted as $\xi_{\mu\nu}(x,\tau)\equiv\xi(x,\tau)g_{\mu\nu}(x,\tau)$. Therefore, the equations of motion take the Langevin form
$\frac{\partial A^{a}_{\mu}(x,\tau)}{{\partial}\tau}=$ $\!\!\!\!\phantom{a}\imath\nabla^{\nu}F^{a}_{\nu\mu}(x,\tau)+\xi^{a}_{\mu}(x,\tau)\,,$ (35)
$\frac{\partial g_{\mu\nu}(x,\tau)}{{\partial}\tau}=$ $\!\!\!\!-\imath\frac{\sqrt{-g}}{2\kappa}\Big{[}R_{\mu\nu}(x,\tau)-\frac{1}{2}R(x,\tau)g_{\mu\nu}(x,\tau)$ (36)
$\!\!\!\!-\kappa T_{\mu\nu}(x,\tau)\Big{]}+\xi_{\mu\nu}(x,\tau)\,.$
Here, the gauge fields and their related strength fields are defined as usual by
$\nabla^{\nu}F^{a}_{\nu\mu}$ $\!\!\!\!=\nabla_{\nu}F^{a\nu}{}_{\mu}$ (37)
$\!\!\!\!=\partial_{\nu}F^{a\nu}_{\mu}+\Gamma^{\nu}_{\nu\rho}F^{a\rho}_{\mu}-\Gamma^{\rho}_{\nu\mu}F^{a\nu}_{\rho}\,,$
$F^{a}_{\mu\nu}$ $\!\!\!\!=\nabla_{\mu}A^{a}_{\nu}-\nabla_{\nu}A^{a}_{\mu}+f^{a}_{\ bc}A^{b}_{\mu}A^{c}_{\nu}$ (38)
$\!\!\!\!=\partial_{\mu}A^{a}_{\nu}-\partial_{\nu}A^{a}_{\mu}+f^{a}_{\ bc}A^{b}_{\mu}A^{c}_{\nu}\,,$
with $f^{abc}$ structure constants of the gauge group, and the geometric quantities fulfil the standard definitions
$\Gamma^{\mu}_{\nu\rho}=\frac{1}{2}g^{\mu\sigma}(\partial_{\nu}g_{\sigma\rho}+\partial_{\rho}g_{\nu\sigma}-\partial_{\sigma}g_{\nu\rho})\,,$ (39)
$R_{\mu\nu}=\partial_{\rho}\Gamma^{\rho}_{\mu\nu}-\partial_{\nu}\Gamma^{\rho}_{\mu\rho}+\Gamma^{\sigma}_{\mu\nu}\Gamma^{\rho}_{\sigma\rho}-\Gamma^{\sigma}_{\mu\rho}\Gamma^{\rho}_{\nu\sigma}\,,$ (40)
with $R\equiv g^{\mu\nu}R_{\mu\nu}$. Finally, the energy momentum tensor of the theory is defined by
$T_{\mu\nu}=\frac{1}{4\pi}F^{a}_{\mu\rho}F_{\nu}{}^{a\rho}-\frac{1}{16\pi}F^{a}_{\rho\sigma}F^{a\ \rho\sigma}g_{\mu\nu}\,,$ (41)
the trace of which, $T\equiv g^{\mu\nu}T_{\mu\nu}$, equates $R$ at equilibrium.
We argue that the out-of-equilibrium Einstein–Yang–Mills system in Eqs. (35)-(36) describes stochastic topology fluctuations in the background manifold, which induces the appearance of center vortices. The appearance of the latter ones, being dual to Wilson loops, is specifically responsible for the asymptotic area-law scaling of Wilson-loop vevs. On the other hand, the appearance of center vortices, as induced by the stochastic gauge-geometry flow, can be interpreted as the emergence of closed chromo-magnetic Faraday lines dual to strings — the chromo-electric fluxes between quarks. Both ’t Hooft loops, responsible for the creation of center vortices, and Wilson loops evolve in the stochastic time $\tau$. The (stochastic) breakdown of the diffeomorphism symmetries entails the generation of topological defects, which can be in turn interpreted as vortices. On the other hand, the stochastic dynamics also induces singularities in the gauge transformation, which are also responsible for the creation of chromo-magnetic vortices as mentioned above. The two stochastic processes are related, and therefore we argue that also the stochastic noise amplitudes for the gravitational and gauge field must be related. It is also relevant to observe that chromo-electric and chromo-magnetic fluxes tangle through an Aharonov-Bohm (geometric) phase. The phase is due to the winding (in space-time) of the chromo-magnetic fluxes around the chromo-electric Wilson loops.
Within our out-of-equilibrium (stochastic) gauge-geometric flow perspective, the winding of Wilson lines can also be described, in a stochastic quantisation approach along the line of Parisi and Wu (1981), as a topological braiding due to the $\theta$-term in QCD. This opens a pathway to a category theory interpretation of confinement, as mainly due to topological effects and topology changes in the stochastic gauge-geometry flow. Consistent with this is our educated guess that the linking number of gluon field lines is at the onset of confinement, being instantiated by stochastic topology transformations of the underlying manifold. When two quarks are moved one away from another, some magnetic closed lines are broken, but then immediately recombine. This induces a breakdown and recombination of the chromo-electric Wilson lines, which are eventually twisted — through the mathematical braiding induced by the Aharonov-Bohm effect — and intertwined to the magnetic flux lines.
Within this framework, we propose that manifolds with the most simple topology (for instance, ×$\mathbb{R}\times S^{3}\cong\mathbb{R}\times S^{3}$) can be conjectured to give rise to ground-states such as the ones of mesons and nucleons. More involved topologies can give rise to multi-quark resonances and other excited meson states.
IV Out-of-equilibrium dynamics and Renormalization Group
The connection between standard RG flow techniques deployed in QFT to investigate the running of gauge and diffeo-invariant observables, and the evolution of the Einstein–Yang–Mills system in a stochastic time deserves a separate analysis that is currently ongoing. Nonetheless, a few interesting features can be immediately derived that concern the appearance in the out-of-equilibrium phase of torsional degrees of freedom, possibly with relevant phenomenological consequences.
What one can see in the case of gauge theories, already in the original paper of Parisi–Wu Parisi and Wu (1981), is that the divergences arising from the singular gauge transformation seem to be translated directly to divergences in the stochastic-time dependence. These divergences are then similar to the secular terms in the multi-scale expansions of differential equations. If we choose the “extended” connection in the projective form
$\bar{\Gamma}^{\alpha}_{\beta\gamma}=\Gamma^{\alpha}_{\beta\gamma}+\delta^{\alpha}_{\gamma}U_{\beta}\,,$ (42)
we immediately derive a conformal transformation of the metric tensor that is sourced by the stochastic flow:
$\frac{\text{d}}{\text{d}\tau}g_{\mu\nu}\left(\tau\right)=-2n^{\alpha}(\tau)U_{\alpha}(\tau)g_{\mu\nu}(\tau)=-2\lambda\left(\tau\right)g_{\mu\nu}\left(\tau\right)\,.$ (43)
This choice of the extended connection is indeed consistent with a rescaling of the metric as a function of the stochastic time and the space-time coordinates, with a scaling factor provided by $\lambda(\tau)=n^{\mu}(\tau)U_{\mu}(\tau)$, which is a local variable.
We may then ask ourselves whether there exists a relation between the scale factor $\lambda$ of the RF and the scale factor of the RG flow. For sure, the value of the scale factor in QFT is not a field itself, i.e. it does not depend on the coordinates. Instead, it would coincide in the case of the Minkowski metric, which is homogeneous and isotropic, with $\lambda$ turning out to depend on $\tau$ only. In a sense we would be “deparametrising” the scale, in a similar way as one deparametrises time in General Relativity.
Another relevant point concerns how to recover the (deparametrised) RG flow of QFT in the low-curvature limit. This was addressed in Ref. Lulli et al. (2023), in light of its relevance for the foundational issue of the collapse of the wave function. This latter was investigated through the lenses of the SRF in Ref. Lulli et al. (2023), and recognised to be induced by the back-reaction with the geometry.
Finally, we mention results that will appear in Lulli et al. (2024), concerning the role of the torsional degrees of freedom in stochastic geometric flows. Specifically, it is possible to show that perturbations of the metric that are out-of-equilibrium introduce a projective torsional term to the connection. Out-of-equilibrium, torsion is dynamical, and therefore the degrees of freedom that are related to it will have to evolve according to the Langevin equation. Asymptotically, the relaxation toward equilibrium that enables to obtain General Relativity is achieved through a corresponding evaporation of the torsional degrees of freedom. Thus, torsional degrees of freedom can be interpreted as a byproduct of the stochasticity of the underlined geometry, and are unfrozen in the out-of-equilibrium phase. Consistently with this picture, the extended connection induces a scale transformation on the reference metric tensor generated by its proper-time derivative.
V Wilsonian approach to QCD and topology changes
We can consider the close neighborhood around a Wilson loop, i.e. a solid torus embedded in $M$ with $\gamma$ being the meridian. We can then consider the knot complement $C(\gamma)=S^{3}-\gamma\times D^{2}$. In a dynamical picture, the knot must fluctuate due to the stochastic RF. For small fluctuations, the knot does not change and it is topologically stable. However, for large fluctuations, two parts of the closed curve come close to each other so that the two parts meet and rearrange in a different way. This operation is well-known in knot theory: one gets a new knot which is concordant to the previous knot. There is a result that every knot is changed by the concordance relation to a hyperbolic knot, i.e. a knot where the knot complement admits a hyperbolic geometry.
The scenario in focus is now the following. At low energies one has small fluctuations and the Wilson loop is determined by the closed curve (i.e. a knot) and the loop is determined by the length of the curve. In contrast, for large fluctuations, one has a change of the knot and then one has to consider to corresponding change of the knot induced by the cobordism between the two knots. Then the Wilson loop is determined by the surface (i.e. the cobordism). This behaviour leads to the confinement (note a Wilson’s original argument about the surface behaviour of the Wilson loop). The size of this surface is confined by the facts that every knot is concordant to a hyperbolic knot and the size of surfaces in hyperbolic manifolds (the knot complement) is determined by Mostow rigidity.
It is possible to interpret stochastic term as a stochastic diffeomorphism, i.e. we consider a function $f:M\rightarrow\mathbb{R}$ on the 3-manifold and the RF:
$\frac{\mbox{d}}{\mbox{d}\tau}g_{\mu\nu}=-2R_{\mu\nu}-2\partial_{\mu}\partial_{\nu}f\,,$ (44)
where $f$ is considered as a generator of diffeomorphisms. This equation can be written as the gradient flow of the functional
[…]
where $f$ can then be interpreted as dilaton.
Usually, a stochastic term is only a continuous function but every homeomorphism of the 3-manifold can be arbitrarily accurate approximate by a diffeomorphism (topological and smooth 3-manifolds are equal). Alternatively, one can also consider a $C^{2}$ function so that $\partial_{\mu}\partial_{\nu}f$ is only continuous.
As a second point, we have to consider the (potential) convergence behaviour of the RF. The behaviour depends strongly on the topology of the 3-manifold and its underlying geometric structure. In Appendix A we describe the eight possible geometries on 3-manifold and the decomposition of the 3-manifold into geometric pieces. For now, we will state the general behaviour: every spherical geometry has a finite extinction form of the RF; furthermore, the other geometries except of the hyperbolic geometry leading to a pinching of the underlying 3-manifold. Therefore, only the RF of the hyperbolic geometry can be written as a stochastic equation away from the fixed point.
VI Proposal for QCD confinement
Geometric fluctuations due to the stochastic RF induce changes in the field lines following the pattern of knot concordance described in Section V. In this setting, the expectation of the knotted and linked Wilson loops is computed by considering the vev of the product of two Wilson loops as
$\langle\mathcal{W}_{\gamma_{1}}(A)\mathcal{W}_{\gamma_{2}}(A)\rangle=$ $\frac{1}{Z}\int\mathcal{D}A\,\mathcal{W}_{\gamma_{1}}(A)\,\mathcal{W}_{\gamma_{2}}(A)$ (46)
[…]
where $S_{\rm YM}$ indicates the Yang-Mills action, and the normalization constant $Z$ is the partition function. Following the analogy of the Chern-Simons case in Cotta-Ramusino et al. (1990), we expect that $\langle\mathcal{W}_{\gamma_{1}}(A)\mathcal{W}_{\gamma_{2}}(A)\rangle$ depends on the linking of $\gamma_{1}$ and $\gamma_{2}$. For Chern-Simons, it is known that switching crossings produces certain skein relations, as shown in Ref. Cotta-Ramusino et al. (1990), which is the reason why the vev of Wilson loops gives link invariants such as the Jones polynomial. In the case of Yang-Mills, we do not expect the result to produce link invariants, but we conjecture that the linking results in interactions that influence confinement. Since linking is a property induced by stochastic fluctuations, as suggested in this article following the stochastic RF, it follows that confinement properties are induced by geometric properties. While we defer a detailed study of Eq. (46) and its relation to confinement to subsequent studies, we emphasize that one can compute $\langle W_{\gamma_{1}}(A)W_{\gamma_{2}}(A)\rangle$ also in the case where $\gamma_{2}$ is obtained from $\gamma_{1}$ by a small shift as in the case of the self-linking number. This corresponds to the presence of self-interaction terms.
VII Conclusions and outlook
We outlined the proposal for a novel geometric phase approach to confinement in Yang-Mills theories. The crucial aspects of the framework we have developed can be summarised as follows:
  • i) hadrons’ ground-states are interpreted geometrically as manifolds. They can be cut away from the vacuum space manifold, since they represent its weakly-interacting sub-manifolds;
  • ii) the gravitational back-reaction to the Yang-Mills fields induces stochastic fluctuations of topology of the sub-manifolds that represent matter;
  • iii) in dual picture, changes of topology corresponds to the generation of gauge vortices, which finally source the geometric phase of hadrons’ ground states.
This mechanism ultimately provides the area law for quarks’ confinement according to the center vortex mechanism. Nonetheless, several relevant points still need to be addressed.
First of all, the formation at the fermi scale of a non-vanishing condensate is eminently related to the topological properties of the manifold. But the energy scales that defines the size of the ground states is still related to dimensional transmutation. Nonetheless, it might be tempting in this framework to relate the energy scale of confinement to amplitude of the stochastic noises that enter the Einstein–Yang-Mills stochastic flow. In Lulli et al. (2022) the amplitude of the gravitational stochastic noise was traced back to the variance of its fluctuations, and shown to be proportional to cosmological constant. On the other hand, for the gauge fields it might be natural to relate the origin of the amplitude of the stochastic noise to chaos effects in the dynamics of the chromo-electric and chromo-magnetic fluxes.
A fascinating possibility, which nonetheless has still to be inspected in detail, is that the stochastic noise of gauge fields could be still decomposed in terms of the gravitational stochastic noise, and finally related to the unbalance of the energy densities of the chromo-electric and chromo-magnetic condensates, as inspected in Addazi et al. (2019a) at the cosmological level. Thus, within this picture, the energy scale of the condensate would be still related to the variance of the stochastic fluctuations of the metric tensor. Outside the fermi world, this acquires the value of the cosmological constant Lulli et al. (2022), but when distributions of energy density are taken into account inside the fermi scale, its value would increase from the one proper of the cosmological constant to that one of $\Lambda_{\rm QCD}$.
Dimensional transmutation can be also reconsidered in the framework we are proposing. We may indeed account for the breakdown of the conformal symmetry at the quantum level, as being induced by the stochastic geometric flow, while the system is out of equilibrium.
The stochastic time regulates the out-of-equilibrium dynamics, making evident that the breakdown of the conformal symmetry (at the quantum level) emerges at non-zero stochastic time distances, and out of equilibrium. Scales can be then individuated in the theory through the coupling between gravity and matter. It is indeed the Newton constant $G$ that provides a dimension-full constant to introduce new physical scale. On the other hand, it is the interplay between the coupling constant of the theory of gravity and the specific coupling constant of matter that generate the physical relevant scale for the QFT under investigation.
Two specific examples of emergence of physical scales through dimensional transmutation – induced by the coupling with matter, and reminiscent of the conformal anomaly as sourced by the trace of the energy-momentum tensor – are provided by the theory of QED (the Bohr radius of the electron, that regulates the “size” of the Hydrogen atom) and the theory of QCD (the $\Lambda_{\rm QCD}$ scale). In these two cases, the physical scales should be provided by a running, induced by the stochastic geometric flow, of the values of $G$, together with a rescaling with the respect to $g_{\rm em}$ and $g_{\rm s}$ of the gravitational coupling constant $G$. Hence, it would be tempting to consider
$r_{\rm Bohr}^{2}\sim G/g_{\rm em}\,,\qquad\delta^{\rm QCD}\Lambda\sim g_{\rm s}/G\,.$ (47)
for the contribution from the QCD sector $\delta^{\rm QCD}\Lambda$ to the cosmological constant $\Lambda$.
The so-called “Vacuum Catastrophe” reflects the fundamental fact that the energy density of the quark-gluon condensate responsible for the color confinement in QCD is off by over forty orders of magnitude and has a wrong sign compared the observable cosmological $\Lambda$-term value. In the Yang-Mills effective action approach Savvidy (1977); Matinyan and Savvidy (1978) it is possible to show that the exact compensation of the electric (metastable) and magnetic (stable) gluon condensate components localised into vacuum domains spatially separated by domain walls in the QCD vacuum (averaged over space-time volumes above the fermi scale) is the necessary and sufficient condition for confinement in QCD Addazi et al. (2019a, b). Provided that the energy scales of “electric gluon” and “magnetic gluon” condensation are not the same, the latter are formed at different space-time separations. However, they evolve towards the same values of the energy density, but with opposite signs, due to their cosmological attractor nature. This effectively causes the cancellation of “electric gluon” and “magnetic gluon” contributions to the QCD ground state in macroscopic volumes, i.e. in the deep infrared limit of the theory Pasechnik et al. (2013a); Addazi et al. (2022). The gravitational interactions then break the discrete mirror symmetry of the QCD vacuum already in the semi-classical approximation. Such a breaking causes a non-zeroth (positive) gravitational correction to the resulting effective QCD-induced cosmological constant $\delta^{\rm QCD}\Lambda$ found to be close (within an order of magnitude) to the observed value Pasechnik et al. (2013b).
Such a cancellation emerges in experiment as a complete disappearance of the colored degrees of freedom from the theory in the IR limit, i.e. with characteristic wave lengths beyond the fermi scale, as they are “trapped” within the domains at the QCD length scale. This is fully compatible with the classical limit of the Yang-Mills theory where the conformal anomaly dynamically vanishes and where only a matter-like (non-relativistic) QCD medium (hadron gas) remains at large separations. For the exact cancellation and, hence, for color confinement the restoration of a discrete (mirror) symmetry between the “electric gluon” and “magnetic gluon” contributions at the level of the effective Lagrangian is an intrinsic property of pure Yang-Mills theory and the RG flow equations.
At the next stage, it would be mandatory to interconnect between the effective Yang-Mills action approach, featuring the dimensional transmutation and an emergence of the mirror symmetry of the Yang-Mills vacua, the respective RG flow and the RF formalism.
Appendix A $3$-Manifold topology and geometry
We hereby collect some general facts on $3$-manifolds and their geometry. A connected 3-manifold $N$ is prime if it cannot be obtained as a connected sum of two manifolds $N_{1}\#N_{2}$ neither of which is the 3-sphere $S^{3}$ (or, equivalently, neither of which is the homeomorphic to $N$). Examples are the 3-torus $T^{3}$ and $S^{1}\times S^{2}$ but also the Poincaré sphere. According to Milnor (1962), any compact, oriented 3-manifold is the connected sum of a unique (up to homeomorphism) collection of prime 3-manifolds (prime decomposition). A subset of prime manifolds are the irreducible 3-manifolds. A connected 3-manifold is irreducible if every differentiable submanifold $S$ homeomorphic to a sphere $S^{2}$ bounds a subset $D$ (i.e. $\partial D=S$) which is homeomorphic to the closed ball $D^{3}$. Moreover, it is known that the only (orientable) prime but reducible 3-manifold is $S^{1}\times S^{2}$.
To understand the geometric properties, a finer decomposition induced by incompressible tori is needed. A properly embedded connected surface $S\subset N$ is called 2-sided if its normal bundle is trivial, and 1-sided if its normal bundle is nontrivial. A 2-sided connected surface $S$ other than $S^{2}$ or $D^{2}$ is called incompressible if for each disk $D\subset N$ with $D\cap S=\partial D$ there is a disk $D^{\prime}\subset S$ with $\partial D^{\prime}=\partial D$. The boundary of a 3-manifold is an incompressible surface. Most importantly, the 3-sphere $S^{3}$, $S^{2}\times S^{1}$ and the 3-manifolds $S^{3}/\Gamma$ with $\Gamma\subset SO(4)$ a finite subgroup do not contain incompressible surfaces. The class of 3-manifolds $S^{3}/\Gamma$ (the spherical 3-manifolds) include cases like the Poincaré sphere ($\Gamma=I^{*}$ the binary icosahedral group) or lens spaces ($\Gamma=\mathbb{Z}_{p}$ the cyclic group).
Let $K_{i}$ be irreducible 3-manifolds containing incompressible surfaces, and assume that $N$ splits into pieces (along embedded $S^{2}$)
[…]
where #$\#_{n}$ denotes the $n$-fold connected sum and $\Gamma\subset SO(4)$ is a finite subgroup. The decomposition of $N$ is unique up to the order of the factors. The irreducible 3-manifolds $K_{1},\ldots,\,K_{n_{1}}$ are able to contain incompressible tori and one can split $K_{i}$ along the tori into simpler pieces $K=H\cup_{T^{2}}G$ Jaco and Shalen (1979) (called the JSJ decomposition). The two classes $G$ and $H$ are the graph manifold $G$ and the hyperbolic 3-manifold $H$.
The hyperbolic 3-manifold $H$ has a torus boundary $T^{2}=\partial H$, i.e. $H$ admits a hyperbolic structure in the interior only. It is also useful to consider the splitting of the link/knot complement. As shown in Budney (2005), the Whitehead double of a knot leads to JSJ decomposition of the complement into the knot complement and the complement of the Whitehead link (along one torus boundary of the Whitehead link complement).
One property of hyperbolic 3-manifolds is central for our considerations: the Mostow rigidity. As shown by Mostow Mostow (1968), every hyperbolic $n-$manifold $n>2$ with finite volume has the property that Every diffeomorphism (especially, every conformal transformation) is induced by an isometry. Therefore, one cannot scale a hyperbolic 3-manifold and the volume is a topological invariant. Together with the prime and JSJ decomposition,
[…]
we can discuss the geometric properties central to Thurstons geometrization theorem: every oriented closed prime 3-manifold can be cut along tori (JSJ decomposition), so that the interior of each of the resulting manifolds has a geometric structure with finite volume. Now, we have to clarify the term ‘geometric structure’. A model geometry is a simply connected smooth manifold $X$ together with a transitive action of a Lie group $G$ on $X$ with compact stabilizers.
A geometric structure on a manifold $N$ is a diffeomorphism from $N$ to $X/\Gamma$ for some model geometry $X$, where $\Gamma$ is a discrete subgroup of $G$ acting freely on $X$. It is a surprising fact that there are also a finite number of three-dimensional model geometries, i.e. only eight different geometries with the following structure: spherical $(S^{3},O_{4}(\mathbb{R}))$, Euclidean $(\mathbb{E}^{3},O_{3}(\mathbb{R})\ltimes\mathbb{R}^{3})$, hyperbolic $(\mathbb{H}^{3},O_{1,3}(\mathbb{R})^{+})$, mixed spherical-Euclidean $(S^{2}\times\mathbb{R},O_{3}(\mathbb{R})\times\mathbb{R}\times\mathbb{Z}_{2})$, mixed hyperbolic-Euclidean $(\mathbb{H}^{2}\times\mathbb{R},O_{1,3}(\mathbb{R})^{+}\times\mathbb{R}\times\mathbb{Z}_{2})$ and 3 exceptional cases called $\tilde{SL}_{2}$ (twisted version of $\mathbb{H}^{2}\times\mathbb{R}$), NIL (geometry of the Heisenberg group as twisted version of $\mathbb{E}^{3}$), SOL (split extension of $\mathbb{R}^{2}$ by $\mathbb{R}$, i.e. the Lie algebra of the group of isometries of 2-dimensional Minkowski space).